OpenStudy (anonymous):

A speed trap is set up with two pressure-activated strips placed across a highway, 130 m apart. A car is speeding along at 33 m/s, while the speed limit is 21 m/s. At the instant the car activates the first strip, the driver begins slowing down. What minimum deceleration is needed in order that the average speed is within the limit by the time the car crosses the second marker?

OpenStudy (anonymous):

yahoo are you actually trying to answer this problem?

OpenStudy (anonymous):

Vavg = Total Displacement/T T = 130/21 = 6.1 s

OpenStudy (anonymous):

s = ut +1/2 a t^2

OpenStudy (anonymous):

now just find out a..))

OpenStudy (anonymous):

hope..thhis helps..

OpenStudy (anonymous):

i love you

OpenStudy (anonymous):

u understood..))

OpenStudy (anonymous):

lol...)) try..to do if u have doubts call me

OpenStudy (anonymous):

wait why choose 21 for the velocity? and isnt there a change of velocity which would make using that formula incorrect?

OpenStudy (anonymous):

u see average speed is within the limit by the time the car crosses the second marker?

OpenStudy (anonymous):

i took it as avg speed

OpenStudy (anonymous):

oh. kk then